OpenStudy (anonymous):

Find the value of x so that the line passing through (10, 5) and (x, 9) has a slope of -2. A. -2 B. 8 C. -8 D. 2

jimthompson5910 (jim_thompson5910):

m = (y2 - y1)/(x2 - x1) -2 = (9 - 5)/(x - 10) -2 = 4/(x - 10) -2(x-10) = 4 ... ... ... x = ???

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome